Find 2 consecutive intergers such that the first less twice the second is equal to 98. My daughter missed this problem on a quiz, and I am trying to help figure it out, so we will understand how to do it next time. Thank You
Answer by nyc_function(2741)

The 2 consecutive integers are x, (x + 1).
2(x + 1) - x = 98
2x + 2 - x = 98
x + 2 = 98
x = 98 - 2
x = 96
The 2 consecutive integers are 96 and 97.
Understand?